TinydinoUwU:

If $1,800 is deposited into an account earning 6% interest, how much will be in the account at the end of 18 years if the interest is compounded weekly?

59mandatory:

5,890.37 in 18 years

TinydinoUwU:

ight thank you

toga:

The formula for compound interest is A = P(1 + r/n)^(nt), where A is the ending amount, P is the principal or initial amount, r is the interest rate, n is the number of times the interest is compounded per year, and t is the time in years. Using the given values, we have: P = 1800 r = 6% = 0.06 n = 52 (since interest is compounded weekly) t = 18 Plugging these values into the formula, we get: A = 1800(1 + 0.06/52)^(52*18) Simplifying this expression, we get: A = 1800(1.001154)^936 A = 1800(3.172) A = 5710.40 Therefore, the amount in the account at the end of 18 years with weekly compounding will be approximately $5,710.40.